Fermi National Laboratory


Accelerator Update

Wednesday June 16
The day shift began with Operations monitoring store 3576 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

By 9:54 AM, a Booster RF station (BRF12) had tripped twice on a "Driver Power Fault." It reset okay. It tripped off again at 10:32 AM.

Operations established beam to the Meson MTest experiment at 11:27 AM and a minute later sent beam to MCenter.

At 11:47 AM, BRF12 tripped again. Operators left the station off and called the RF technicians.

MCenter finished using their beam at 12:53 PM.

Pbar experts began a target scan at 1:04 PM.

A Main Injector (MI) expert began a coalescing study at 1:05 PM.

A MI expert began a barrier bucket study at 2:19 PM.

A TeV expert changed the separator helix voltage. This opened the helix a little less than 5%.

Operations halted stacking to prepare Pbar for a reverse proton study at 6:57 PM.

Operations resumed sending beam to MCenter at 11:23 PM.

Operations reported at 11:58 PM that BRF12 had tripped off several times.

Reverse proton studies ended and Pbar resumed stacking at 11:58 PM.

Thursday June 17
The midnight shift began with Operations monitoring store 3576, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

A Tevatron (TeV) expert began an end-of-store study at 4:18 AM. This study was to determine the amount of beam in the beam abort gap.

Operations terminated store 3576 at 5:05 AM.

Operators made the Collider Detector at Fermilab (CDF) and D-zero (D0) collision halls and TeV sectors A-E ready for access by experimenters and techs checking for vacuum leaks at 5:19 AM.

MCenter finished using beam at 8:26 AM.

Operators prepared Linac for an access at 10:05 AM. This access lasted about forty-five minutes.

At 10:58 AM, cryo system techs reported that they were changing a TeV sector A1 wet engine flywheel. At 11:22 AM, cryo system techs reported that they had to either rebuild or replace the wet engine for TeV sector F2.

Operators prepared Booster for an hour access at 11 AM.

Operations reported at 2:05 PM that Pbar had been having problems since 9:45 AM with bad emittances and stacking losses. Operators and experts continued to investigate.

Operations reported at 2:54 PM that Booster's East Anode power supply (EAPS) had tripped off again; it had tripped off six times since in the last hour.

At 3 PM, Operations reported that Pbar seemed to have fixed itself. The investigation continued.

BRF12 tripped off at 3:22 PM. It reset okay.

Both D0 and CDF were out of access by 4 PM.

The Booster EAPS tripped off at 4:52 PM. Operations could find nothing wrong and turned the supply back on.

Operations established beam to MTest at 5:46 PM, and to MCenter by 8:40 PM.

TeV experts began orbit studies at 10:11 PM.

Friday June 18
The midnight shift began with Operations monitoring the Recycler stash, with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE and MTest taking beam.

Operations established beam to MCenter at 1:12 AM.

At 1:54 AM, Operations reported concern over Pbar's emittance blow up. Operators contacted Pbar experts who suggested some tune changes. These changes helped.

At 2:53 AM, Operations had trouble with MI coalescing. Upon investigation operators discovered three MI RF stations off, in local, and with their alarms bypassed. Operators turned the stations on.

Operations loaded a 12x0 proton only study-store in the TeV at 3 AM.

Operations put the TeV into shot setup at 5:03 AM. Switchyard (SY) beam gets disabled during shot setup.

Operations established store 3582 at 7:13 AM with an initial luminosity of 75.04E30. Operators immediately began working on the colliding beam problems with collimator moves for CDF.

BRF12 tripped off at 7:26 AM. It reset okay.

Pbar resumed stacking at 7:54 AM.

Operations resumed sending beam to the Meson experiments.

Plans for Today
The plans for today and this weekend are to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/18/2004   email Fermilab